OpenStudy (anonymous):

If there are initially 2000 bacterial in a culture, and the number of bacteria doubles each hour, the number of bacteria after t hours can be found using the formula N = 2000(2t). How many bacteria will be present after 7 hours?

OpenStudy (anonymous):

Okay, they give you the formula. t = hours N = 2000(2(7)) N = 2000(14) N = 28000

OpenStudy (anonymous):

Got it? :P

OpenStudy (anonymous):

ooohh no!

OpenStudy (anonymous):

??

OpenStudy (anonymous):

\[ N(t) = 2000\times 2^t\] doubles every hour compute \[N(7)=2000\times 2^7\]

OpenStudy (anonymous):

thanks satellite73

OpenStudy (anonymous):

yw